OpenStudy (anonymous):

HELP plz! The probability tht an integrated circuit chip will have defective etching is 0.06, the probability tht it will have acrack defect is 0.03, and the probability tht it has both defects is 0.02 (a) what is the probability tht a newly manufactured chip will have either an etching or a crack defect? (b) what is the probability tht a newly manufactured chip will have neither defect?

OpenStudy (amistre64):



OpenStudy (amistre64):

P(AorB) = P(A) +P(B) -P(AandB) P(-A and -B) = 1 - P(AorB)

OpenStudy (anonymous):

Thanks amistre64 :)

OpenStudy (anonymous):

Amistre64 If the qts is tht fint the probability tht it will get one of the two defects dn???